EXPLAIN THE FORMULA  lim x→0 (sin x0 ) /x  = Π / 1800

Grade:12th Pass

1 Answers

RAJORSHI PAUL
12 years ago

Note that the result lim (x->0)sin x/x =1 holds only when x is in radians (& not in degrees ) (You can check the derivation of the result)

So x*(x degree) =x(in radian).(180*/pi) So sin x* =sin(180*x/pi)

So lim(x->0) sin x*/x = lim(x->0) sin (180*x/pi)/x = pi/180*